Ball A rolls off a speed of 1.2m/s,east and strikes a second stationary ball B.Ball B is twice as a massive as ball A.If after collision,ball A is observed to be travelling north at a speed of 0.4 m/s,find the velocity of Ball B after collision? With complete solution


1
Expert's answer
2020-11-30T14:52:02-0500
"\bold{P_A=P_A'+P_B'}"

"2mv_x'=mv_x\\v_x'=0.5v_x=0.5(1.2)=0.6\frac{m}{s}\\v_y'=0.5v_y=0.5(0.4)=0.2\frac{m}{s}"

"v'=\sqrt{0.6^2+0.2^2}=0.63\frac{m}{s}"

"\theta=\arctan{\frac{0.2}{0.6}}=18.4\degree\text{ S of E}"